# Comprehensive Benchmarking Environment for Worker Flexibility in Flexible Job Shop Scheduling Problems

### v1

@@ 3 Representations of the Flexible Job Shop Scheduling Problem
This section establishes the foundation for setting up the benchmarking environment by describing three different modeling approaches for the FJSSP. The first of these is a Mixed Integer Linear Programming (MILP) representation defined in Sec. 3.1. It allows exact methods of mathematical programming to be applied. Furthermore, an implementation as Constraint Programming (CP) model is provided in Sec. 3.2. Toolboxes for this type of problem representation are currently among the most sophisticated solving techniques for Production Scheduling tasks. Lastly, a convenient problem representation for Simulation-Based Optimization (SBO) algorithms, e.g. meta-heuristics, is presented in Sec. 3.3. It includes the definition of a precise schedule encoding and the design of the production system as an functional simulation environment.
3.1 Mixed-Integer Linear Programming Formulation
@@
The Mixed Integer Linear Program (MILP) formulation is a proven model for the FJSSP to address these problem class using mathematical optimization methods, i.e. branch-and-bound or cutting-plane methods. In this section, the MILP is systematically formulated describing the decision variables, constraints and objective functions in detail.
Accordingly, the decision variables are defined in the following way:
@@
3.1.1 General restrictions in FJSSP
@@
3.1.2 Objective function and associated constraints
@@
3.2 Constraint Programming Formulation
The FJSSP can be easily formulated as a discrete finite constraint optimization problem(COP). The two COP solvers that we take into account in this comparison implement variable types, like interval variables and constraints like no-overlap or end-before-start, that are tailored to deal with scheduling problems.
In this comparison OR-Tools (ORT) an open-source solver developed by Google and CP-Optimizer (CPO) a constraint optimization solver provided as a part of the IBM ILOG CPLEX Optimization Studio are used. The problem formulation for both solvers uses the same subscripts, variables, and index sets as the MILP formulation (see 3.1).
3.2.1 Variables
@@ 3.2.3 Constraints
3.3 Simulation-Based Optimization Attempt
@@
4 Expansion of the FJSSP for Worker Flexibility
The Flexible Job Shop Scheduling Problem with Worker Flexibility (FJSSP-W) is an extension to the FJSSP which adds worker flexibility. This means that for every operation, an eligible worker also needs to be assigned to operate the machine to which the operation is assigned. The selected machine-worker combination further influences the processing time of the operation on the machine. In addition to the FJSSP, the FJSSP-W therefore introduces a set of workers $\mathcal{W}=\{W_{1},\dots,W_{o}\}$. Each operation $O_{i,j}$ needs to be assigned to a permissible machine $M_{k}\in\mathcal{M}_{i,j}\subset\mathcal{M}$ and a worker $W_{s}\in\mathcal{W}_{i,j,k}\subset\mathcal{W}$ the set of permissible workers to process the operation $O_{i,j}$ on the assigned machine $M_{k}$. Similar to the FJSSP, the FJSSP-W assumes that all operations, machines, and workers are operational and available at time 0. The objective for the FJSSP-W is to find those operation-machine-worker combination that minimize the makespan $C$. As opposed to the FJSSP, the flexibility $\beta$ of a benchmark instance needs to take more into account than only the machines. All possible assignment combination for every operation need to be taken into account. This is the same as for the FJSSP, but includes the possible worker assignments for each operation-machine combination as well. The average amount of assignment options $\omega_{avg}$ is determined by Eq. (20).
$\omega_{all}=\sum_{i=1}^{\texttt{N}}(|\mathcal{M}_{i}|+\sum_{j=1}^{|\mathcal{M}_{i}|}|\mathcal{W}_{\mathcal{M}_{i},j}|)$ (19)
$\omega_{avg}=\frac{1}{\texttt{N}}\sum_{i=1}^{n}\sum_{j=1}^{n_{i}}\sum_{k=1}^{m}|\mathcal{W}_{ijk}|\quad$ (20)
Which leads to $\beta$ defined as Eq. (21), with $\omega_{uni}$ being the overall amount of unique machine-worker combinations available for the benchmark instance.
$\beta=\cfrac{\omega_{avg}}{\omega_{uni}}$ (21)
In the same vein, the definition for $dv$ for FJSSP-W needs to be rewritten as Eq. (22)
$\begin{split}dv=\cfrac{d_{unique}}{\sum_{i=1}^{\texttt{N}}(|\mathcal{M}_{i}|+\sum_{j=1}^{|\mathcal{M}_{i}|}|\mathcal{W}_{\mathcal{M}_{i},j}|)}\end{split}$ (22)
4.1 Design Proposal of Worker Flexibility Integration
The FJSSP-W considers that operations in the production process need workers to be present. Therefore, workers also need to be assigned to the different operations. In the FJSSP-W, it is considered that not all workers will work at the same speed, and not all workers are capable of performing every operation with ever (or any) of the available machines for the operation. To receive a set of benchmark problems to test and develop appropriate approaches, the given FJSSP benchmark instances are extended to contain the additional required information about the possible worker assignments. To that end, we provide an algorithm similar to [20] , which is capable of extending any FJSSP instance in the format of the gathered benchmark instances (Alg. 1). The parameters used for the benchmark generation are the amount of workers (default: $1.5*m$), the upper and lower bound range for the durations (default: $ub=1.1$, $lb=0.9$ of the original duration) to have a random fluctuation in the durations required by the workers of ±$\pm 10\%$, and optionally a random seed for reproducibility.
After the initialization of the required parameters, in line 1 an empty list for saving the result benchmark is created. In line 2, the original benchmark file, located at $path$, is read and saved into a list, separated into the rows of the file. The base information about the benchmarks are added to the result, as a list representing a new row in the output, in line 3, keeping the same format as the original files, with the added information about the amount of workers available for the problem instance. Starting from line 4 and ranging to line 25, the rest of the rows of the original benchmark file are processed one by one. In line 5, a variable to keep track of the reading position in the current row ($idx$) is initialized, as well as an empty list, which is added to the output as a new row. The amount of operations needed for the job represented by the row is added to the new row in line 6. From line 7 to line 24, the current row from the original benchmark, according to the original format of the problem instances, is processed. To do so, $idx$ is raised by $1$ in row 8 to address the next operation that needs to be parsed. In line 9, the amount of machines available for the operation is read, and added to the output row in line 10. From line 11 to line 23, the different machine assignments and their respective durations for the current operation are parsed. To achieve that, line 12 raises $idx$ by $1$ again to advance the reading position of the input current row. Since $idx$ is now referencing to a machine ID in the original benchmark, this possible machine assignment is also added to the output row in line 13. In line 14, the amount of workers which are going to be able to perform the required operation on this specific machine is chosen uniformly from the interval [$1$, $n_{workers}$]. The amount of workers available for this machine assignment ($workers$) is added to the output row in line 15. In line 16, the available workers for the assignment are selected randomly, using a function named $Sample$, using the parameters $n_{workers}$ and $workers$ to select the IDs of the workers in the interval [0, $n_{workers}$). Note that generally the assigned IDs of machines and workers are converted to 0-based indexing for easier implementation as opposed to the original 1-based indexing. The subset of workers is stored in the list $\mathbf{selected}$. To advance the reading position again, $idx$ is raised by $1$ in line 17. $idx$ now points at the original duration of the machine assignment for the given operation, which is stored in $d_{orig}$ in line 18. Starting from line 19 and ending in line 22, the current operation-machine assignments are finalized by choosing the durations for the different worker assignments. To do this, the duration for each worker $d_{work}$ is uniformly chosen from the interval [$d_{orig}*lb$, $d_{orig}*ub$] and rounded to the closest integer in line 20. Line 21 adds the worker and the duration for the operation-machine-worker combination to the output row in $out$.
1:$n_{workers}>0$, $lb\geq 0$, $ub\geq 0$, $path$
2:$\mathbf{out}\leftarrow\{\}$
3:$\mathbf{rows}\leftarrow ReadFile(path)$
[…]
5:for $i\leftarrow 1$ to $||\mathbf{rows}||$ do
6:     $\mathbf{out}_{i}\leftarrow\{\}$, $idx\leftarrow 0$
7:     $\mathbf{out}_{i}\cup\mathbf{rows}_{i,idx}$
8:     for $j\leftarrow 1$ to $||\mathbf{rows}_{i}||$ do
9:         $idx\leftarrow idx+1$
10:         $m\leftarrow\mathbf{rows}_{i,idx}$
11:         $\mathbf{out}_{i}\cup\{m\}$
12:         for $k\leftarrow 1$ to $m$ do
13:              $idx\leftarrow idx+1$
14:              $\mathbf{out}_{i}\cup{\mathbf{rows}_{i,idx}}$
15:              $workers\leftarrow Uniform(1,n_{workers})$
16:              $\mathbf{out}_{i}\cup\{workers\}$
17:              $\mathbf{selected}\leftarrow Sample(n_{workers},workers$)
18:              $idx\leftarrow idx+1$
19:              $d_{orig}\leftarrow\mathbf{rows}_{i,idx}$
20:              for $l\leftarrow 1$ to $workers$ do
21:                  $d_{work}\leftarrow Round(Uniform(d_{orig}*lb,d_{orig}*ub))$
22:                  $\mathbf{out}_{i}\cup\{selected_{l},d_{work}\}$
23:              end for
24:         end for
25:     end for
26:end for
After this routine is finished, the benchmark is extended with worker flexibility and kept similar to the original format. The original format can be seen in an arbitrarily constructed example in Listing 2.
12 3
23 2 0 10 2 15 3 0 10 1 5 2 10 2 0 10 2 15
32 1 1 15 2 1 10 2 15
Listing 2: Original Benchmark example
Row 1 of Listing 2 contains the information that the problem instance consists of 2 jobs and 3 available machines. In row 2 it is specified that Job 1 consists of 3 operations. Operation 1 of Job 1 can be processed on 2 machines (Machine 0 with processing time 10 and Machine 1 with processing time 15). Operation 2 of Job 1 can be processed on 3 machines. Either Machine 0 with processing time 10, Machine 1 with processing time 5 or Machine 2 with processing time 10. Operation 3 of Job 1 has 2 machines available (Machine 0 with processing time 10 and Machine 2 with processing time 15). Job 2 requires 2 operations. Operation 1 of Job 2 can be processed on only 1 machine (Machine 1 with processing time 15), while Operation 2 of Job 2 can be processed on 2 machines (Machine 1 with processing time 10 and Machine 2 with processing time 15).
In the new format of the benchmarks, line 1 contains the general information about the instance (amount of jobs, amount of machines, amount of workers), followed by a row by row description of every job that needs to be scheduled. Every row starts with the number of operations required for the job to be finished. Following this first row entry, every operation starts with the amount of machines required, followed by the entries for every machine. Every entry for the machine starts with the machine id, followed by the amount of workers available to process the operation on this specific machine. After that, for every worker available on the machine, an entry consisting of the worker id, followed by the duration of the operation-machine-worker combination is listed in the row.
An arbitrarily constructed example for this format can be seen in Listing 3. REPLACE:
1 2 3 4
2 3 2 0 2 1 10 3 5 2 1 0 15 3 0 1 2 10 1 2 0 10 3 5 2 2 2 5 3 15 2 0 2 2 15 3 10 2 2 0 15 1 5
3 2 1 1 3 1 10 2 5 3 10 2 1 1 2 10 2 2 0 10 3 15
Listing 3: Benchmark with worker flexibility example
This example means that, according to row 1, there are 2 jobs, 3 machines and 4 workers in this production environment. Job 1 (row 2) consists of 3 operations. Operation 1 can be processed on 2 machines. The first of these machines is Machine 0, which has 2 workers available to perform this operation. Worker 1 with a processing time 10 and Worker 3 with a processing time of 5. The second machine to process Operation 1 of Job 1 is Machine 2. Machine 2 has 1 worker available (Worker 0) with a processing time 15. Operation 2 of Job 1 has 3 machines available. Machine 0, with 1 worker (Worker 2) with processing time 10, Machine 1 with 2 workers (Worker 0 with processing time 10 and Worker 3 with processing time 5), and Machine 2 with 2 workers (Worker 2 with processing time 5 and Worker 3 with processing time 15). Operation 3 of Job 1 can be processed on 2 machines. Machine 0, which has 2 possible worker assignments (Worker 2 with processing time 15 and Worker 3 with processing time 10), and Machine 2. Machine 2 also has 2 possible workers (Worker 0 with processing time 15 and Worker 1 with processing time 5). Job 2 (row 3), requires 2 operations to be processed. Operation 1 of Job 2 has 1 possible machine (Machine 1) with 3 possible workers. Worker 1 with processing time 10, Worker 2 with processing time 5 and Worker 3 with processing time 10. Operation 2 of Job 2 can be assigned to 2 machines. Either Machine 1, which has 1 possible worker (Worker 2 with processing time 10), or Machine 2, with the possibility to assign the operation to 2 different workers (Worker 0 with processing time 10 or Worker 3 with processing time 15).
4.2 Mixed-Integer Linear Programming Formulation
An obvious generalization of the MILP formulation for the FJSSP (3.1) is used. The worker flexibility is modeled by an additional set of constraints that follows the same ideas as the sequencing constraints on machine level in 3.1. The following description explains in detail only the additional notations and constraints of the MILP formulation for the FJSSP-W.
4.2.1 Indices and Variables
The decision variables $Y_{i,j,k}$ in 3.1 have to be changed to $Y_{i,j,k,s}$ which are 1, if operation $O_{i,j}$ is processed on machine $M_{k}$ by worker $W_{s}$ and 0 else. Furthermore an additional type of decision variables is necessary: $U_{i,j,i^{\prime},j^{\prime}}$, which is set to 1 if Operation $O_{i,j}$ is processed after operation $O_{i^{\prime},j^{\prime}}$ by the same worker. To describe the admissible workers for processing operation $O_{i,j}$ on machine $M_{k}$ the index set
is used.
4.2.2 Constraints
Assignment constraints
Sequencing constraints on job level
Sequencing constraints on machine level
Sequencing constraints on worker level
As in 4.2.2 these two sets of constraints have to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ and operation $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$. But contrary to 4.2.2 they have to hold for machines $M_{k}$ ($k\in M_{i,j}$) and $M_{k^{\prime}}$ ($k^{\prime}\in M_{i^{\prime},j^{\prime}}$) and for workers $W_{s}$ with $s\in W_{i,j,k}\cap W_{i^{\prime},j^{\prime},k^{\prime}}$. For the large number $L$, one can use
The idea of these constraints is the same as in Eq. (9) and (10) explained there in detail. Finally makespan and variable type constraint are the same as in 3.1.
4.3 Constraint Programming Formulation
for $1\leq i\leq n$ and $1\leq j\leq n_{i}$.
The sequencing constraints on job level are the same as in 3.2:
and the sequencing constraints on machine level are very similar to 3.2. For $1\leq k\leq m$ the following constraints are added:
4.4 Simulation-Based Optimization Attempt
The extension to the encoding for simulation-based approaches stays the same for sequence and machine assignments as in section 3.3. To add the worker flexibility, an additional vector is added. The vector has the same properties as the machine assignment vector, in that it is in the same fixed order and contains the IDs of the workers assigned to the operation-machine combination chosen to do the task. Extending the example used in section 3.3 with 3 available workers for the system, a problem instance could look like Listing 4.
@@ 4.4 Simulation-Based Optimization Attempt
Listing 4: Example sequence, assignment and worker vector

### v2

@@ Acknowledgements
The following sections describe three types of problem formulations. The formulations are provided for the FJSSP instances, and for the FJSSP-W instances, respectively. In Sec. A, a Mixed Integer Linear Programming (MILP) representation is presented that allows for the application of exact Mathematical Programming solvers like Gurobi. Further, an implementation as Constraint Programming (CP) model is provided in Sec. B. Toolboxes (e.g. CPLEX) for this type of problem representation are currently among the most sophisticated solving techniques for Production Scheduling tasks. Finally, a convenient problem representation for Simulation-Based Optimization (SBO) algorithms (e.g. meta-heuristics, like GA) is presented in Sec C. It includes the definition of a precise schedule encoding and the design of the production system as an functional simulation environment.
Appendix A Mixed-Integer Linear Programming Formulation
@@
The Mixed Integer Linear Program (MILP) formulation is a common model to approach the FJSSP(-W) problem class using mathematical optimization methods, i.e. branch-and-bound or cutting-plane methods. In this section, the MILP representation is systematically formulated describing the decision variables, constraints and objective functions in detail. All necessary parameters are first specified for the FJSSP before an extension for the FJSSP-W is provided.
A.1 Decision Variables of the FJSSP
In a first step, the decision variables can be defined in the following way:
@@ A.1 Decision Variables of the FJSSP
A.2 General Restrictions in FJSSP
@@
A.3 Objective function and associated constraints for the FJSSP
@@
A.4 Extended MILP Formulation for the FJSSP-W
An obvious generalization of the MILP formulation for the FJSSP (LABEL:sec:fjssp-milp) is used. The worker flexibility is modeled by an additional set of constraints that follows the same ideas as the sequencing constraints on machine level in LABEL:sec:fjssp-milp. The following description explains in detail only the additional notations and constraints of the MILP formulation for the FJSSP-W.
A.4.1 Indices and Variables
The decision variables $Y_{i,j,k}$ in LABEL:sec:fjssp-milp have to be changed to $Y_{i,j,k,s}$ which are 1, if operation $O_{i,j}$ is processed on machine $M_{k}$ by worker $W_{s}$ and 0 else. Furthermore an additional type of decision variables is necessary: $U_{i,j,i^{\prime},j^{\prime}}$, which is set to 1 if Operation $O_{i,j}$ is processed after operation $O_{i^{\prime},j^{\prime}}$ by the same worker. To describe the admissible workers for processing operation $O_{i,j}$ on machine $M_{k}$ the index set
is used.
A.4.2 Constraints
Assignment constraints
Sequencing constraints on job level
Sequencing constraints on machine level
Sequencing constraints on worker level
As in A.4.2 these two sets of constraints have to hold for operations $O_{i,j}$ with $1\leq i<n$ and $1\leq j\leq n_{i}$ and operation $O_{i^{\prime},j^{\prime}}$ with $i^{\prime}>i$ and $1\leq j^{\prime}\leq n_{i^{\prime}}$. But contrary to A.4.2 they have to hold for machines $M_{k}$ ($k\in M_{i,j}$) and $M_{k^{\prime}}$ ($k^{\prime}\in M_{i^{\prime},j^{\prime}}$) and for workers $W_{s}$ with $s\in W_{i,j,k}\cap W_{i^{\prime},j^{\prime},k^{\prime}}$. For the large number $L$, one can use
The idea of these constraints is the same as in Eq. (14) and (15) explained there in detail. Finally makespan and variable type constraint are the same as in A.3.
Appendix B Constraint Programming Formulation
The FJSSP can be easily formulated as a discrete finite single-objective Constraint Optimization Problem (COP). The two Constraint Programming (CP) solvers that we take into account in this comparison implement variable types, like interval variables and constraints like no-overlap or end-before-start, that are tailored to deal with scheduling problems of general form.
The investigations included the solvers OR-Tools (ORT), i.e. an open-source solver developed by Google, and CP-Optimizer (CPO), a constraint optimization solver provided as a part of the IBM ILOG CPLEX Optimization Studio. The problem formulation for both solvers uses the same subscripts, variables, and index sets as the MILP formulation (see A).
B.1 Variables
@@ B.3 Constraints
B.4 Extending the CP Formulation for the FJSSP-W
for $1\leq i\leq n$ and $1\leq j\leq n_{i}$.
The sequencing constraints on job level are the same as in LABEL:sec:fjssp-cp:
and the sequencing constraints on machine level are very similar to LABEL:sec:fjssp-cp. For $1\leq k\leq m$ the following constraints are added:
Appendix C Simulation-Based Optimization Attempt
@@
C.1 Extension of SBO Encoding for FJSSP-W
The extension to the encoding for simulation-based approaches stays the same for sequence and machine assignments as in section LABEL:sec:fjssp-sbo. To add the worker flexibility, an additional vector is added. The vector has the same properties as the machine assignment vector, in that it is in the same fixed order and contains the IDs of the workers assigned to the operation-machine combination chosen to do the task. Extending the example used in section LABEL:sec:fjssp-sbo with 3 available workers for the system, a problem instance could look like Listing 2.
@@ C.1 Extension of SBO Encoding for FJSSP-W
Listing 2: Example sequence, assignment and worker vector